# Factor-Driven Network Informed Restricted Vector Autoregression

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 6th ACM International Conference on AI in Finance; November 15–18, 2025; Singaporebooktitle: 6th ACM International Conference on AI in Finance (ICAIF ’25), November 15–18, 2025, Singapore
@@ 1.4. Notation
[…]
@@ 3. Estimation
Algorithm 1 NIRVAR estimation of the idiosyncratic component.
@@
2:Compute the sample covariance matrix $\hat{\Gamma}_{\xi}\leftarrow\hat{\xi}\hat{\xi}^{\prime}\!/\!T$.
3:Obtain the $d$ eigenvectors $\bm{v}_{1},\dots,\bm{v}_{d}$ corresponding to the $d$ largest eigenvalues in magnitude of $\hat{\Gamma}_{\xi}$, and form the $N\times d$ matrix $V$ with columns $\bm{v}_{1},\dots,\bm{v}_{d}$.
4:Cluster the rows of $V$ via a $K$-component Gaussian mixture model.
5:Set $\hat{z}(i)$ to the cluster assignment of row $i$ for each $i\in[N]$.
6:Define $\hat{A}_{ij}=\mathds{1}\{\hat{z}(i)=\hat{z}(j)\}$.
7:Estimate the parameters $\tilde{\Phi}$ of the restricted VAR model $\bm{\xi}_{t}=\hat{A}\odot\tilde{\Phi}\bm{\xi}_{t-1}+\bm{\epsilon}^{(\hat{A})}_{t}$ via OLS.
8:Return Cluster labels $\hat{z}(i)$; estimated coefficient matrix $\hat{\Phi}$.
@@ 4.3. Violating the positive eigengap assumption
We simulate data from the FNIRVAR model for different values of $N$, with $T=1000$, and monitor the growth of the eigenvalues of $\hat{\Gamma}$ as $N$ increases. For this simulation, we choose a single factor ($r=1$) and five blocks ($K=5$). The loadings $\Lambda$ are sampled from a mixture of normals: $\Lambda_{i,j}\sim 0.5\cdot\mathcal{N}(1,\sigma_{\Lambda}^{2})+0.5\cdot\mathcal{N}(-1,\sigma_{\Lambda}^{2})$. The factor VAR coefficient matrices are sampled independently from $\mathcal{N}(0,0.1)$ and then rescaled such that $\rho(P^{*})=0.6$ to ensure stationarity. The number of common shocks is chosen to be equal to the number of factors ($q=r$) and $\Gamma_{u}$ is set to the identity matrix. The number of factor lags is chosen to be $l_{F}=5$. The block matrix $B$ is set to have diagonal values of 0.950.950.95 and off-diagonal values of 0.010.010.01. The weights for the NIRVAR coefficient matrix are chosen as $\tilde{\Phi}_{ij}\sim\mathcal{N}(\mu_{i},1)$ where $\mu_{i}\coloneqq(-1)^{z(i)+1}z(i)$, then symmetrized as $\tilde{\Phi}\leftarrow(\tilde{\Phi}+\tilde{\Phi}^{\prime})/2$. The NIRVAR coefficient matrix $\Phi=A\odot\tilde{\Phi}$ is then normalised such that $\rho(\Phi)=0.9$. The idiosyncratic error covariance matrix is chosen to be the identity: $\Gamma_{\epsilon}=I_{N}$. For each $N$, 20 replicate datasets are simulated using the above hyperparameters. We control the eigengap $\Delta$ through the choice of $\sigma_{\Lambda}$. For $\sigma_{\Lambda}^{2}=9\times 10^{-3}$, $\Delta>0$ whereas for $\sigma_{\Lambda}^{2}=2.5\times 10^{-4}$, $\Delta<0$. Figure 1 shows the minimum eigenvalue of $\hat{\Gamma}_{\chi}$ and maximum eigenvalue of $\hat{\Gamma}_{\xi}$ as a function of $N$ for the two scenarios: $\Delta>0$ and $\Delta<0$. In both cases, the eigenvalue of $\hat{\Gamma}_{\chi}$ grows linearly in $N$, as is expected for a strong factor (see Barigozzi and Hallin, 2024, for further discussion on the growth rate of eigenvalues of factor models). However, when $\Delta<0$ (the case where Assumption 1 is violated), $\lambda_{\text{max}}(\hat{\Gamma}_{\chi})<\lambda_{\text{min}}(\hat{\Gamma}_{\xi})$, and therefore, removing the top eigenvalues of $\hat{\Gamma}$, as in our two-step estimation procedure, will not pick out the common component. As a further observation, we note that the eigenvalues of $\hat{\Gamma}$ do not grow linearly with $N$, confirming that the idiosyncratic component of FNIRVAR cannot be interpreted as a strong factor model.
@@
Figure 1. Growth of eigenvalues of $\hat{\Gamma}$ with $N$ under a FNIRVAR model when Assumption 1 (a) holds and (b) is violated.
@@ 5.1. Application to daily returns prediction
The task is to predict the next day pvCLCL market excess returns. We adopt a rolling window backtesting framework with a fixed four year look-back. The predictive performance between 01/01/2004 and 31/12/2020 of the three models are compared via commonly used metrics in the financial literature (Gu et al., 2020). The mean number of factors chosen by PCp2 (Bai and Ng, 2002) over the backtest period was 12.1612.1612.16, whilst the mean order of the factor VAR model was 3.973.973.97.
@@
We also consider decile portfolios, in which solely the top $x$largest in magnitude signals are traded each day (Fama and French, 1992). The Sharpe ratio and mean daily PnL (using FNIRVAR as the prediction model) for equal weighted decile portfolios with different transaction costs are shown in Table 4. The table shows that lower decile portfolio have higher PnL, with the Sharpe ratios maintaining a similar level.
@@
A value-weighted portfolio accounts for the liquidity in each asset and is defined by $\omega_{i}^{(t)}=\min(\alpha\nu_{i,t},\beta),$ where $\nu_{i,t}=\text{median}\{V_{i,t},\dots,V_{i,1}\},$ with $V_{i,t}$ being the dollar volume traded for asset $i$ on day $t$. The parameters $\alpha$ and $\beta$ control the imbalance between portfolio weights: as $\alpha$ increases and $\beta$ decreases we get closer to the equal weighted portfolio setting. Table 5 shows that as $\alpha$ increases and $\beta$ decreases, the Sharpe ratio increases and portfolio variance decreases, respectively. The mean portfolio weight for every choice of $(\alpha,\beta)$ was 0.00154. We also show the total book value and mean bet size in dollars for different values of $(\alpha,\beta)$ in Table 6. We note that the largest book size considered is \$156.9\\$156.9\$ 156.9 million which is much smaller that the total market. Therefore, market impact will be negligible. For fixed $\alpha=0.001$ and $\beta=500,000$, Table 7 shows the Sharpe ratio and mean daily PnL for various decile portfolios and different transaction costs. Note that the choice of $(\alpha,\beta)$ in Table 7 is the furthest from an equal-weighted portfolio out of all pairs of $(\alpha,\beta)$ considered. Again, lower decile portfolios give larger PnL.
@@
Metric Total bet size (×$\times 10^{6}$) Mean bet size per asset (×$\times 10^{3}$)
@@ 6. Conclusion and discussion
In this work, we introduced a factor-driven network-based VAR model called FNIRVAR for modelling the common and idiosyncratic components of a high dimensional multivariate time series. In FNIRVAR, a static factor model is combined with a network VAR model whose coefficient matrix is the weighted adjacency matrix of a SBM. The regularisation due to the network makes FNIRVAR an appropriate model for high dimensional settings. Under the assumption of strong factors and a large eigengap, we estimate the factors and loadings via PCA. NIRVAR estimation is employed on the remaining idiosyncratic component leading to a block-sparse estimated VAR coefficient matrix. As the FNIRVAR estimator does not require an observed network, it is especially suited to applications in finance and economics where human-defined networks may not be available. To test the predictive performance of FNIRVAR, we constructed a backtesting framework for three datasets of daily returns, intraday returns, and macroeconomic variables. In all three cases FNIRVAR out-performed a static factor model as well as a static factor plus LASSO estimated sparse VAR model. FNIRVAR showed particularly compelling results on the financial applications, achieving a Sharpe ratio of 1.951.951.95 for the task of daily returns prediction, and 2.202.202.20 for the task of thirty minutely intraday returns prediction. These results highlight the benefits of including sparse estimation along with dense factor modelling.

### v2

@@ Abstract.
conference: 6th ACM International Conference on AI in Finance; November 15–18, 2025; Singapore, Singaporebooktitle: 6th ACM International Conference on AI in Finance (ICAIF ’25), November 15–18, 2025, Singapore, Singaporedoi: 10.1145/3768292.3770412isbn: 979-8-4007-2220-2/2025/11
@@ 1.3. Related literature
Finally, we mention the growing use of modern machine learning techniques for time series prediction tasks. Gu et al. (2020) provide a comprehensive overview and comparison of machine learning methods for empirical asset pricing, and suggest that the inclusion of nonlinear predictor interactions can lead to predictive gains missed by more traditional regression-based strategies.
@@ 1.4. Notation
For an integer $k$, we let $[k]$ denote the set $\{1,\dots,k\}$. The $n\times n$ identity matrix is written $I_{n}$, and the indicator function is $\mathds{1}\{\mathcal{B}\}=1$ if the event $\mathcal{B}$ occurs and 0 otherwise. For vectors, $\bm{v}_{1},\dots,\bm{v}_{n}\in\mathbb{R}^{p}$, we let $M=(\bm{v}_{1},\dots,\bm{v}_{n})\in\mathbb{R}^{p\times n}$ be the matrix whose columns are given by $\bm{v}_{1},\dots,\bm{v}_{n}$. The transpose of $M$ is denoted $M^{\prime}$. We write $M_{i,:}$ and $M_{:,j}$ to denote the $i$-th row and $j$-th column of $M$ considered as vectors, respectively. We use $\odot$ to denote the Hadamard product between two matrices of the same dimensions. For matrices $M_{1},\dots,M_{r}$, $(M_{1}|\cdots|M_{r})$ and $(M_{1};\dots;M_{r})$ denote the column-wise and row-wise concatenation of the matrices, respectively. We write $\mathcal{S}_{K-1}=\{x\in\mathbb{R}^{K}:\sum_{k=1}^{K}x_{k}=1,x_{k}>0,k\in[K]\}$ to denote the interior of the unit simplex in $K$ dimensions.
@@ 3. Estimation
Algorithm 1 FNIRVAR estimation.
@@
2:Estimate $r$ factors and corresponding loadings via static PCA: $\widehat{\Lambda}=E$ and $\widehat{F}_{t}=E^{\prime}\bm{x}_{t}$ where $E:=\text{eig}(XX^{\prime}/T)_{[:,1:r]}$.
3:Estimate $l_{F}$ factor VAR coefficient matrices $P_{k}$ via least-squares.
4:Estimate the idiosyncratic component as $\hat{\bm{\xi}}_{t}=\bm{X}_{t}-\hat{\bm{\chi}}_{t}$.
5:Compute the sample covariance matrix $\hat{\Gamma}_{\xi}\leftarrow\hat{\xi}\hat{\xi}^{\prime}\!/\!T$.
6:Obtain the $d$ eigenvectors $\bm{v}_{1},\dots,\bm{v}_{d}$ corresponding to the $d$ largest eigenvalues in magnitude of $\hat{\Gamma}_{\xi}$, and form the $N\times d$ matrix $V$ with columns $\bm{v}_{1},\dots,\bm{v}_{d}$.
7:Cluster the rows of $V$ via a $K$-component Gaussian mixture model.
8:Set $\hat{z}(i)$ to the cluster assignment of row $i$ for each $i\in[N]$.
9:Define $\hat{A}_{ij}=\mathds{1}\{\hat{z}(i)=\hat{z}(j)\}$.
10:Estimate the parameters $\tilde{\Phi}$ of the restricted VAR model $\bm{\xi}_{t}=\hat{A}\odot\tilde{\Phi}\bm{\xi}_{t-1}+\bm{\epsilon}^{(\hat{A})}_{t}$ via OLS.
11:Return Factors $\widehat{F}_{t}$, loadings $\widehat{\Lambda}$, factor VAR coefficient matrices $\widehat{P}_{k},$ cluster labels $\hat{z}(i)$; estimated coefficient matrix $\hat{\Phi}$.
@@
The computational complexity of FNIRVAR is dominated by estimation of the idiosyncratic component, which is $O(N^{3}/K^{2}+TN^{2}/K)$ when $\Gamma_{\epsilon}$ is homoskedastic and $O(N^{3}+TN^{2})$ otherwise (since generalised least-squares estimation must be used in this case) (Martin et al., 2024). Estimation of the factors and loadings via PCA on the other hand can be done using a Lanczos type algorithm in $O(TNr)$ time. Least-squares estimation of the factor VAR process is $O(Tr^{2}l_{F}^{2})$ and is therefore subdominant since $r\ll N$ and $l_{F}\ll N$. We note that least-squares estimation of an unrestricted VAR is $O(N^{3}+TN^{2})$, and thus, FNIRVAR estimation is faster whenever $\Gamma_{\epsilon}$ is assumed to be homoskedastic.
@@ 4.3. Violating the positive eigengap assumption
We simulate data from the FNIRVAR model for different values of $N$, with $T=1000$, and monitor the growth of the eigenvalues of $\hat{\Gamma}$ as $N$ increases. For this simulation, we choose a single factor ($r=1$) and five blocks ($K=5$). The loadings $\Lambda$ are sampled from a mixture of normals: $\Lambda_{i,j}\sim 0.5\cdot\mathcal{N}(1,\sigma_{\Lambda}^{2})+0.5\cdot\mathcal{N}(-1,\sigma_{\Lambda}^{2})$. The factor VAR coefficient matrices are sampled independently from $\mathcal{N}(0,0.1)$ and then rescaled such that $\rho(P^{*})=0.6$ to ensure stationarity. The number of common shocks is chosen to be equal to the number of factors ($q=r$) and $\Gamma_{u}$ is set to the identity matrix. The number of factor lags is chosen to be $l_{F}=5$. The block matrix $B$ is set to have diagonal values of $0.95$ and off-diagonal values of $0.01$. The weights for the NIRVAR coefficient matrix are chosen as $\tilde{\Phi}_{ij}\sim\mathcal{N}(\mu_{i},1)$ where $\mu_{i}\coloneqq(-1)^{z(i)+1}z(i)$, then symmetrized as $\tilde{\Phi}\leftarrow(\tilde{\Phi}+\tilde{\Phi}^{\prime})/2$. The NIRVAR coefficient matrix $\Phi=A\odot\tilde{\Phi}$ is then normalised such that $\rho(\Phi)=0.9$. The idiosyncratic error covariance matrix is chosen to be the identity: $\Gamma_{\epsilon}=I_{N}$. For each $N$, 20 replicate datasets are simulated using the above hyperparameters. We control the eigengap $\Delta$ through the choice of $\sigma_{\Lambda}$. For $\sigma_{\Lambda}^{2}=9\times 10^{-3}$, $\Delta>0$ whereas for $\sigma_{\Lambda}^{2}=2.5\times 10^{-4}$, $\Delta<0$. Figure 1 shows the minimum eigenvalue of $\hat{\Gamma}_{\chi}$ and maximum eigenvalue of $\hat{\Gamma}_{\xi}$ as a function of $N$ for the two scenarios: $\Delta>0$ and $\Delta<0$. In both cases, the eigenvalue of $\hat{\Gamma}_{\chi}$ grows linearly in $N$, as is expected for a strong factor (see Barigozzi and Hallin, 2024, for further discussion on the growth rate of eigenvalues of factor models). However, when $\Delta<0$ (the case where Assumption 1 is violated), $\lambda_{\text{max}}(\hat{\Gamma}_{\chi})<\lambda_{\text{min}}(\hat{\Gamma}_{\xi})$, and therefore, removing the top eigenvalues of $\hat{\Gamma}$, as in our two-step estimation procedure, will not pick out the common component. As a further observation, we note that the eigenvalues of $\hat{\Gamma}$ do not grow linearly with $N$, confirming that the idiosyncratic component of FNIRVAR cannot be interpreted as a strong factor model.
@@
Figure 1. Growth of eigenvalues of $\hat{\Gamma}$ with NN under a FNIRVAR model when Assumption 1 (a) holds and (b) is violated.
@@ 5.1. Application to daily returns prediction
The task is to predict the next day pvCLCL market excess returns. We adopt a rolling window backtesting framework with a fixed four year look-back. The predictive performance between 01/01/2004 and 31/12/2020 of the three models are compared via commonly used metrics in the financial literature (Gu et al., 2020). The mean number of factors chosen by PCp2 (Bai and Ng, 2002) over the backtest period was $12.16$, whilst the mean order of the factor VAR model was $3.97$.
@@
We also consider decile portfolios, in which solely the top xx% largest in magnitude signals are traded each day (Fama and French, 1992). The Sharpe ratio and mean daily PnL (using FNIRVAR as the prediction model) for equal weighted decile portfolios with different transaction costs are shown in Table 4. The table shows that lower decile portfolio have higher PnL, with the Sharpe ratios maintaining a similar level.
@@
A value-weighted portfolio accounts for the liquidity in each asset and is defined by $\omega_{i}^{(t)}=\min(\alpha\nu_{i,t},\beta),$ where $\nu_{i,t}=\text{median}\{V_{i,t},\dots,V_{i,1}\},$ with $V_{i,t}$ being the dollar volume traded for asset $i$ on day $t$. The parameters $\alpha$ and $\beta$ control the imbalance between portfolio weights: as $\alpha$ increases and $\beta$ decreases we get closer to the equal weighted portfolio setting. Table 5 shows that as $\alpha$ increases and $\beta$ decreases, the Sharpe ratio increases and portfolio variance decreases, respectively. The mean portfolio weight for every choice of $(\alpha,\beta)$ was 0.00154. We also show the total book value and mean bet size in dollars for different values of $(\alpha,\beta)$ in Table 6. We note that the largest book size considered is $\mathdollar 156.9$ million which is much smaller that the total market. Therefore, market impact will be negligible. For fixed $\alpha=0.001$ and $\beta=500,000$, Table 7 shows the Sharpe ratio and mean daily PnL for various decile portfolios and different transaction costs. Note that the choice of $(\alpha,\beta)$ in Table 7 is the furthest from an equal-weighted portfolio out of all pairs of $(\alpha,\beta)$ considered. Again, lower decile portfolios give larger PnL.
@@
Metric Total bet size ($\times 10^{6}$) Mean bet size per asset ($\times 10^{3}$)
@@ 6. Conclusion and discussion
In this work, we introduced a factor-driven network-based VAR model called FNIRVAR for modelling the common and idiosyncratic components of a high dimensional multivariate time series. In FNIRVAR, a static factor model is combined with a network VAR model whose coefficient matrix is the weighted adjacency matrix of a SBM. The regularisation due to the network makes FNIRVAR an appropriate model for high dimensional settings. Under the assumption of strong factors and a large eigengap, we estimate the factors and loadings via PCA. NIRVAR estimation is employed on the remaining idiosyncratic component leading to a block-sparse estimated VAR coefficient matrix. As the FNIRVAR estimator does not require an observed network, it is especially suited to applications in finance and economics where human-defined networks may not be available. To test the predictive performance of FNIRVAR, we constructed a backtesting framework for three datasets of daily returns, intraday returns, and macroeconomic variables. In all three cases FNIRVAR out-performed a static factor model as well as a static factor plus LASSO estimated sparse VAR model. FNIRVAR showed particularly compelling results on the financial applications, achieving a Sharpe ratio of $1.95$ for the task of daily returns prediction, and $2.20$ for the task of thirty minutely intraday returns prediction. These results highlight the benefits of including sparse estimation along with dense factor modelling.